# PCoreSet: Effective Active Learning through Knowledge Distillation from Vision-Language Models

### v1

@@ Appendix C Summary of Datasets
Caltech101 [20] 100 4,128 1,649 2,465 General objects
DTD [15] 47 2,820 1,128 1,692 Textures
EuroSAT [33] 10 13,500 5,400 8,100 Satellite
FGVCAircraft [57] 100 3,334 3,333 3,333 Aircraft
Food101 [9] 101 50,500 20,200 30,300 Food
Flowers102 [63] 102 4,093 1,633 2,463 Plants
OxfordPets [65] 37 2,944 736 3,669 Animals
StanfordCars [47] 196 6,509 1,635 8,041 Vehicles
SUN397 [91] 397 15,880 3,970 19,850 Scenes
UCF101 [80] 101 7,639 1,898 3,783 Actions
ImageNet [73] 1,000 1.28M - 50,000 General objects
We provide the details of datasets we used in our experiments in this section. Our experimental evaluation encompasses 11 diverse datasets, including ImageNet [73] and 10 additional datasets spanning various domains and classification challenges. These datasets represent a broad spectrum of visual recognition tasks including fine-grained classification across multiple domains: generic object recognition [20], automobile classification [47], flower species identification [63], aircraft categorization [57], pet breed classification [65], food recognition [9], scene understanding [91], texture classification [15], satellite imagery analysis [33], and human action recognition [80]. Details of number of samples in each training, validation and test splits are illustrated in the Table 3.
We adhere to the train, validation, and test splits established in prior work [41]. Our experimental protocol begins with a 1-shot setting, where only a single labeled example per class is available, with all remaining images in the training set treated as unlabeled. This approach extends to subsequent active learning rounds, where the unlabeled set consists of all training images not currently in the labeled set. To further simulate realistic constraints where validation data may be inaccessible, particularly in scenarios beginning with minimal labeled examples and progressively acquiring annotations, we conduct our experiments without utilizing the validation sets.
@@
Caltech101 [20] 41.28 16 400 56.81 4,128
DTD [15] 60.00 60 60 0.00 2,820
EuroSAT [33] 1,350.00 1,000 1,500 174.80 13,500
FGVCAircraft [57] 33.34 33 34 0.48 3,334
Food101 [9] 500.00 500 500 0.00 50,500
Flowers102 [63] 40.13 20 129 22.20 4,093
OxfordPets [65] 79.57 74 80 1.26 2,944
StanfordCars [47] 33.21 19 54 3.47 6,509
SUN397 [91] 40.00 40 40 0.00 15,880
UCF101 [80] 75.63 58 97 10.72 7,639
ImageNet [73] 1,281.17 732 1,300 70.22 1,281,167
We further presents class balance statistics for each dataset in the Table 4 to provide additional insight into inherent class imbalances, a phenomenon that several researchers have addressed within active learning selection processes [3, 8, 37]. It is important to note that our research addresses a distinct problem formulation; we primarily investigate teacher model bias rather than dataset bias. Notably, our findings demonstrate that the teacher bias emerges even in balanced datasets, showing the effectiveness of our approach in active knowledge distillation scenario.
@@ D.1 Experiment on MobileNetV2
We further validate our methodology across diverse model architectures, including the lightweight MobileNetV2 [75]. The experimental results are presented in Fig. 13. For these experiments, we maintained the training configuration established for ResNet-18 [29], employing a ResNet-50 CLIP teacher [68] with identical hyperparameters. Our empirical results demonstrate that PCoreSet consistently outperforms alternative selection methods in the active distillation setting, exhibiting performance trends that align with our primary ResNet-18 [29] experiments.
@@ D.2 Active Learning without Knowledge Distillation
The main purpose of our work is to propose utilizing a foundational teacher model directly in the knowledge distillation process [34]. Thus, we conducted active learning experiments without knowledge distillation as our baseline to enable direct comparison. Using ResNet-18 [29] and following our main experimental protocol in Appendix B, we performed comparative analysis of knowledge distillation effects in the active learning process. The results are presented in Fig. 14. Interestingly, widely-adopted baselines such as uncertainty [35], coreset [76], and badge [4] do not consistently outperform random or class-balanced selection methods Alg. 5 on datasets such as FGVC [57], with our PCoreSet Alg. 2 achieving superior performance. This occurs despite the approximately equal distribution of classes in these datasets as shown in Table 4 in the Appendix C, suggesting that maintaining probabilistic balance may be beneficial even in traditional active learning scenarios. However, PCoreSet does not demonstrate clear effectiveness in this traditional active learning setting, indicating that maintaining probabilistic diversity is specifically effective in distillation scenarios, where it serves to leverage teacher’s structured bias propagation.
@@ D.3 Few-shot Teacher Distillation
In real-world applications, it is more natural to consider scenarios where both teacher and student models evolve concurrently. Specifically, we can apply few-shot learning methods to the generalist foundational teacher model with additional labeled samples acquired in each round, though our main experiments focused on zero-shot teachers to validate the effectiveness of integrating knowledge distillation into the active learning framework. Therefore, we extended our experiments to incorporate few-shot teachers that are updated with newly labeled samples in each acquisition round. For this purpose, we employed CLAP [77], a few-shot learning method built upon CLIP that does not require a validation set—an appropriate choice for our setting where we assume extremely limited data availability (e.g., 1-shot training datasets) without access to validation data. The experimental results for student model performance are presented in Fig. 15, with corresponding teacher model performance shown in Fig. 16. In the few-shot teacher scenario, we observed consistent trends where probabilistic diversity effectively leverages bias, resulting in enhanced performance. Moreover, examining the few-shot teacher performance in Fig. 16 reveals that our PCoreSet outperforms alternative selection methods for teacher models as well. We attribute this to the fact that samples selected to leverage teacher bias naturally serve as beneficial examples for improving the teacher model, simultaneously enhancing performance on both the teacher and student sides.
@@ E.1 Heatmap of Selected Samples
We visualize the distribution of selected samples using heatmaps in Fig. 17.
@@ E.2 Qualitative Results on Selected Samples from PCoreSet
We visualize samples selected by PCoreSet from various datasets in Figures 18 through 24, showcasing how PCoreSet selects diverse and representative examples across different visual domains. These visualizations provide qualitative evidence of PCoreSet’s effectiveness in identifying informative samples that help leverage the structured bias of the teacher model.

### v2

@@ Appendix C Summary of Datasets
Caltech101 (Fei-Fei et al., 2004) 100 4,128 1,649 2,465 General objects
DTD (Cimpoi et al., 2014) 47 2,820 1,128 1,692 Textures
EuroSAT (Helber et al., 2019) 10 13,500 5,400 8,100 Satellite
FGVCAircraft (Maji et al., 2013) 100 3,334 3,333 3,333 Aircraft
Food101 (Bossard et al., 2014) 101 50,500 20,200 30,300 Food
Flowers102 (Nilsback & Zisserman, 2008) 102 4,093 1,633 2,463 Plants
OxfordPets (Parkhi et al., 2012) 37 2,944 736 3,669 Animals
StanfordCars (Krause et al., 2013) 196 6,509 1,635 8,041 Vehicles
SUN397 (Xiao et al., 2010) 397 15,880 3,970 19,850 Scenes
UCF101 (Soomro et al., 2012) 101 7,639 1,898 3,783 Actions
ImageNet (Russakovsky et al., 2015) 1,000 1.28M - 50,000 General objects
We provide the details of datasets we used in our experiments in this section. Our experimental evaluation encompasses 11 diverse datasets, including ImageNet (Russakovsky et al., 2015) and 10 additional datasets spanning various domains and classification challenges. These datasets represent a broad spectrum of visual recognition tasks including fine-grained classification across multiple domains: generic object recognition (Fei-Fei et al., 2004), automobile classification (Krause et al., 2013), flower species identification (Nilsback & Zisserman, 2008), aircraft categorization (Maji et al., 2013), pet breed classification (Parkhi et al., 2012), food recognition (Bossard et al., 2014), scene understanding (Xiao et al., 2010), texture classification (Cimpoi et al., 2014), satellite imagery analysis (Helber et al., 2019), and human action recognition (Soomro et al., 2012). Details of number of samples in each training, validation and test splits are illustrated in the Tab.˜3.
We adhere to the train, validation, and test splits established in prior work (Kang et al., 2025). Our experimental protocol begins with a 1-shot setting, where only a single labeled example per class is available, with all remaining images in the training set treated as unlabeled. This approach extends to subsequent active learning rounds, where the unlabeled set consists of all training images not currently in the labeled set. To further simulate realistic constraints where validation data may be inaccessible, particularly in scenarios beginning with minimal labeled examples and progressively acquiring annotations, we conduct our experiments without utilizing the validation sets.
@@
Caltech101 (Fei-Fei et al., 2004) 41.28 16 400 56.81 4,128
DTD (Cimpoi et al., 2014) 60.00 60 60 0.00 2,820
EuroSAT (Helber et al., 2019) 1,350.00 1,000 1,500 174.80 13,500
FGVCAircraft (Maji et al., 2013) 33.34 33 34 0.48 3,334
Food101 (Bossard et al., 2014) 500.00 500 500 0.00 50,500
Flowers102 (Nilsback & Zisserman, 2008) 40.13 20 129 22.20 4,093
OxfordPets (Parkhi et al., 2012) 79.57 74 80 1.26 2,944
StanfordCars (Krause et al., 2013) 33.21 19 54 3.47 6,509
SUN397 (Xiao et al., 2010) 40.00 40 40 0.00 15,880
UCF101 (Soomro et al., 2012) 75.63 58 97 10.72 7,639
ImageNet (Russakovsky et al., 2015) 1,281.17 732 1,300 70.22 1,281,167
We further presents class balance statistics for each dataset in the Tab.˜4 to provide additional insight into inherent class imbalances, a phenomenon that several researchers have addressed within active learning selection processes (Aggarwal et al., 2020; Bengar et al., 2022; Huang et al., 2024). It is important to note that our research addresses a distinct problem formulation; we primarily investigate teacher model bias rather than dataset bias. Notably, our findings demonstrate that the teacher bias emerges even in balanced datasets, showing the effectiveness of our approach in active knowledge distillation scenario.
@@ D.1 Experiment on MobileNetV2
We further validate our methodology across diverse model architectures, including the lightweight MobileNetV2 (Sandler et al., 2018). The experimental results are presented in Fig.˜11. For these experiments, we maintained the training configuration established for ResNet-18 (He et al., 2016), employing a ResNet-50 CLIP teacher (Radford et al., 2021) with identical hyperparameters. Our empirical results demonstrate that PCoreSet consistently outperforms alternative selection methods in the active distillation setting, exhibiting performance trends that align with our primary ResNet-18 (He et al., 2016) experiments.
@@ D.2 Active Learning without Knowledge Distillation
The main purpose of our work is to propose utilizing a foundational teacher model directly in the knowledge distillation process (Hinton et al., 2015). Thus, we conducted active learning experiments without knowledge distillation as our baseline to enable direct comparison. Using ResNet-18 (He et al., 2016) and following our main experimental protocol in Appendix˜B, we performed comparative analysis of knowledge distillation effects in the active learning process. The results are presented in Fig.˜12. Interestingly, widely-adopted baselines such as uncertainty (Holub et al., 2008), coreset (Sener & Savarese, 2017), and badge (Ash et al., 2019) do not consistently outperform random or class-balanced selection methods Algorithm˜5 on datasets such as FGVC (Maji et al., 2013), with our PCoreSet Algorithm˜2 achieving superior performance. This occurs despite the approximately equal distribution of classes in these datasets as shown in Tab.˜4 in the Appendix˜C, suggesting that maintaining probabilistic balance may be beneficial even in traditional active learning scenarios. However, PCoreSet does not demonstrate clear effectiveness in this traditional active learning setting, indicating that maintaining probabilistic diversity is specifically effective in distillation scenarios, where it serves to leverage teacher’s structured bias propagation.
@@ D.3 Few-shot Teacher Distillation
In real-world applications, it is more natural to consider scenarios where both teacher and student models evolve concurrently. Specifically, we can apply few-shot learning methods to the generalist foundational teacher model with additional labeled samples acquired in each round, though our main experiments focused on zero-shot teachers to validate the effectiveness of integrating knowledge distillation into the active learning framework. Therefore, we extended our experiments to incorporate few-shot teachers that are updated with newly labeled samples in each acquisition round. For this purpose, we employed CLAP (Silva-Rodriguez et al., 2024), a few-shot learning method built upon CLIP that does not require a validation set—an appropriate choice for our setting where we assume extremely limited data availability (e.g., 1-shot training datasets) without access to validation data. The experimental results for student model performance are presented in Fig.˜13, with corresponding teacher model performance shown in Fig.˜14. In the few-shot teacher scenario, we observed consistent trends where probabilistic diversity effectively leverages bias, resulting in enhanced performance. Moreover, examining the few-shot teacher performance in Fig.˜14 reveals that our PCoreSet outperforms alternative selection methods for teacher models as well. We attribute this to the fact that samples selected to leverage teacher bias naturally serve as beneficial examples for improving the teacher model, simultaneously enhancing performance on both the teacher and student sides.
D.4 Logit Adjustment for Calibrated Teacher Prediction
In this section, we investigate whether logit adjustment (Menon et al., 2020), a technique designed for imbalanced learning, can mitigate the structured prediction bias in teacher models. While logit adjustment (LA) provides a principled approach to handling class imbalance, we demonstrate that it is orthogonal to our PCoreSet method and can be used in conjunction to potentially enhance performance.
Logit adjustment (Menon et al., 2020) addresses skewed, imbalanced class distributions in training data by applying a principled and effective post-hoc adjustment based on empirical class frequencies. Formally, $l^{\prime}:=l-\tau\log\pi\in\Delta^{C-1}$, where $l$ and $l^{\prime}$ denote the original and adjusted logits, respectively. $\pi\in\Delta^{C-1}$ is the class prior estimated from the empirical distribution, i.e., $\pi:=\mathbb{E}_{x\sim p(x)}[p(y\mid x)]\approx\tfrac{1}{N}\sum_{n=1}^{N}y_{n}$, with $y_{n}\in\{0,1\}^{C}$ one-hot labels in the supervised training dataset $\mathcal{D}^{(l)}$.
Given that teacher predictions in our setting are often skewed and biased, we explore whether adjusting teacher logits can provide an improved inductive bias. Our hypothesis is that PCoreSet can then leverage this calibrated teacher signal more effectively.
We consider two logit-adjustment variants:
  1. 1.
    Hard LA. Let $\mathcal{D}=\mathcal{D}^{(l)}\cup\mathcal{D}^{(u)}$. Estimate the class prior by pseudo-label frequencies on $\mathcal{D}$ with teacher predictions $f(x)$:
    $\pi_{c}\approx\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\mathbf{1}\left[\arg\max_{c^{\prime}}[f(x)]_{c^{\prime}}=c\right],\qquad\pi=[\pi_{1},\ldots,\pi_{C}]^{\top},$ (23)
    where $\mathbf{1}$ is an indicator function.
  2. 2.
    Soft LA. Estimate the class prior by the average teacher probabilities:
    $\pi\approx\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}f(x).$ (24)
We conducted experiments on 8 of the 10 datasets by integrating LA into teacher predictions. Base results are averaged over 5 seeds, while Soft LA results are averaged over 3 seeds.
Table 5: Logit adjustment experiments across different active learning strategies. Numbers in parentheses show the difference from the baseline (no LA).
Strategy LA Active Learning Round
2 3 4 5 6 7 8
Random - 65.1 67.4 69.5 71.3 72.4 73.9 74.5
Hard 55.1(-10.1) 59.9(-7.4) 63.1(-6.5) 66.0(-5.3) 67.5(-5.0) 69.4(-4.5) 70.6(-3.9)
Soft 64.4(-0.7) 67.6(+0.2) 69.9(+0.3) 71.5(+0.2) 72.8(+0.4) 74.0(+0.1) 75.0(+0.5)
Coreset - 65.7 67.7 69.8 71.5 72.4 73.5 74.8
Hard 55.5(-10.2) 60.6(-7.1) 63.6(-6.2) 66.6(-4.9) 68.8(-3.6) 70.1(-3.4) 71.9(-2.9)
Soft 65.3(-0.4) 68.1(+0.4) 70.4(+0.6) 71.4(-0.1) 73.0(+0.6) 73.6(+0.2) 75.3(+0.5)
Uncertainty - 63.8 66.5 67.9 69.8 71.1 72.4 73.4
Hard 52.6(-11.2) 57.4(-9.1) 61.0(-6.9) 63.7(-6.1) 65.6(-5.5) 68.5(-3.9) 69.9(-3.5)
Soft 63.3(-0.5) 66.3(-0.2) 68.1(+0.2) 70.0(+0.2) 71.1(+0.0) 72.3(-0.1) 73.2(-0.2)
BADGE - 65.3 68.1 70.3 72.1 73.3 74.3 75.1
Hard 54.6(-10.7) 59.6(-8.4) 63.7(-6.6) 66.4(-5.7) 68.0(-5.2) 69.9(-4.4) 71.1(-4.0)
Soft 64.5(-0.8) 67.9(-0.2) 70.2(-0.1) 72.2(+0.1) 73.4(+0.1) 74.5(+0.2) 75.6(+0.5)
ClassBalanced - 65.6 68.3 70.8 72.2 73.3 74.1 75.3
Hard 55.4(-10.2) 58.8(-9.4) 62.6(-8.2) 65.7(-6.5) 67.2(-6.1) 69.2(-4.9) 70.3(-5.0)
Soft 65.1(-0.5) 69.1(+0.8) 71.3(+0.5) 72.9(+0.7) 74.1(+0.8) 75.0(+0.9) 75.7(+0.4)
PCoreSet - 65.7 68.6 71.0 72.8 74.1 75.3 76.2
Hard 56.1(-9.5) 61.8(-6.8) 64.6(-6.4) 67.8(-5.1) 70.4(-3.6) 71.7(-3.6) 73.4(-2.8)
Soft 65.9(+0.2) 69.2(+0.6) 71.5(+0.5) 73.1(+0.3) 74.1(+0.1) 75.3(+0.01) 76.3(+0.01)
As detailed in Tab.˜5, we observe that Soft LA slightly improves PCoreSet on average, even without exhaustive tuning of the hyperparameter $\tau$. These results indicate that while logit adjustment can provide marginal benefits, the core strength of PCoreSet lies in its ability to effectively leverage structured teacher prediction bias within the ActiveKD framework, yielding substantial gains across diverse datasets even without additional calibration techniques.
@@ E.1 Heatmap of Selected Samples
We visualize the distribution of selected samples using heatmaps in Fig.˜15.
@@ E.2 Qualitative Results on Selected Samples from PCoreSet
We visualize samples selected by PCoreSet from various datasets in Figures 16 through 22, showcasing how PCoreSet selects diverse and representative examples across different visual domains. These visualizations provide qualitative evidence of PCoreSet’s effectiveness in identifying informative samples that help leverage the structured bias of the teacher model.
@@
Appendix F The Use of LLMs
We used LLMs solely for light editing such as correcting grammatical errors and polishing some words. They did not contribute to research ideation, experiments, analysis, or substantive writing.